OpenStudy (anonymous):

Please help! Medal and Fan! Is –(27 + (–3)) = –27 + 3 A. commutative property B. associative property C. opposite of a sum property

OpenStudy (melissa_something):

Funny I saw something like this earlier haha.. I think its C

OpenStudy (melissa_something):

The property of the opposite of a sum is referring to the fact that the opposite of a sum of real numbers is equal to the sum of the opposites. For instance: for all real numbers a and b we can see that -(a+b) = (-a) + (-b) original soruce_ http://www.ask.com/question/property-of-the-opposite-of-a-sum

OpenStudy (melissa_something):

if not B

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@Melissa_Something Do you know this one? 35 + (–15 + 4) = (35 + (–15)) + 4 A. commutative property B. associative property C. opposite of a sum property